1. Interchange the order of integration and evaluate the integral S S Siny dy dx. 2. Use Lagrange multipliers to find the p question 2
Answer #1

Da f = x 9-16 distance of point (2,9, 2) form angin da bi ry + 2² det egenys) esitytz we have to manimize g(2,9, 2) Og = (22So, 2 cases y (250) d = 4+1+0=5 2) (-2, 0) d = 4+1to 5 So there are 2 points nearest to от – (2,1,0) and (-2,1,0) the.

.

Please give it a thumbs up if this solution helped you. Still any doubt ? Comment box is all yours, I will be happy to help you :)